The ratio of 150cm to 1metre is 1:1.5.
The ratio of 150cm to 1metre is 1:1.5.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The given statement is False
We know that, 1 metre = 100 cm

So, the ratio of 150cm to 1metre is = 150/100 = 15/10

Divide both numerator and denominator by 5 we get, = 3/2

So, the ratio of 150cm to 1metre is 3: 2


Thanks